Find the equation of the line containing the point (4,-2)and having the slope m=6
Answer by alicealc(293)

line equation with the slope of m and passes through (x1, y1) is:
y - y1 = m*(x -x1)

so, if the line has the slope of 6 and passes through (4, -2) then the line equation is:
y - (-2) = 6*(x - 4)
y + 2 = 6x - 24
y = 6x - 24 - 2
y = 6x - 26